Grade 11General Physics

When a thin plano convex lens (n = 1.5) polished from its plane surface behave like a concave mirror of focal length 40 cm. The radius of curvature of lens is

Aryan
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
 
Dear Aryan
Thr focal length of convex side of plano convex length is 40 cm. Because after the plane surface, it will behave like a concave mirror of focal length = f(lens)/2
f(mirror) = 0.4 m
f(lens) = 0.8 m
From lens maker formula
1/f = (\LARGE \mu -1) [1/Rc -  1/Rp]
1/0.8 = 0.5 *1/Rc
Rc = 0.4 m = 40 cm
 
Regards
Arun (askIITians forum expert)